Calculate the electric field associated with a LASER beam having energy density 10^6 J/cm^2
1
Expert's answer
2020-01-10T09:50:12-0500

Energy density is given by formula

"W=\frac{εε_0E×E}{2} (1)"

Using (1) we get

"E=\sqrt{\frac{2w}{εε_0}} (2)"

In our case, w=106 J/cm^2, ε=1, ε0=8.85×10-12

We got

E=47×107 J


Answer

47×107 J